# Influence of thermal noise on the field-driven dynamics of the non-collinear antiferromagnet Mn3Sn

### v1

@@ Acknowledgment
  • Depopulation factor with $H_{\mathrm{ay}}$
The energy of the system, given by the the Hamiltonian at the saddle point$\Big{(}\theta_{\mathrm{oct}}=\pi/2,\varphi_{\mathrm{oct}}=\sin^{-1}(\frac{H_{\mathrm{y,eff}}}{H_{\mathrm{K}}})/\pi-\sin^{-1}(\frac{H_{\mathrm{y,eff}}}{H_{\mathrm{K}}})\Big{)}$, is given as  Rózsa et al. (2019); Konakanchi et al. (2025):
@@
and, therefore, the relationship between $m_{\mathrm{oct},z}$ and $\varphi_{\mathrm{oct}}$ is given as
@@
×$\times\int_{\pi/2}^{\pi}\alpha(-3M_{\mathrm{s}}H_{\mathrm{K}}\sin\varphi_{\mathrm{oct}}\cos\varphi_{\mathrm{oct}}+3M_{\mathrm{s}}H_{\mathrm{y,eff}}\cos\varphi_{\mathrm{oct}})$
×$\times\sqrt{\frac{H_{\mathrm{J}}H_{\mathrm{K}}}{H_{\mathrm{K}}^{2}\sin^{2}\varphi_{\mathrm{oct}}+H_{\mathrm{y,eff}}^{2}-2H_{\mathrm{K}}H_{\mathrm{y,eff}}\sin\varphi_{\mathrm{oct}}}}\quantity(2H_{\mathrm{K}}^{2}\sin\varphi_{\mathrm{oct}}\cos\varphi_{\mathrm{oct}}-2H_{\mathrm{K}}H_{\mathrm{y,eff}}\cos\varphi_{\mathrm{oct}})d\varphi_{\mathrm{oct}}.$
Numerically integrating the above equation enables us to compute the inverse of the depopulation factor $\quantity[A\quantity(\frac{\delta E}{k_{\mathrm{B}}T})]^{-1}$. For the range of $H_{\mathrm{ay}}$ and $\mathcal{V}$ considered in this work, $\quantity[A\quantity(\frac{\delta E}{k_{\mathrm{B}}T})]^{-1}\approx 1$ with less than $5\%$ error, as presented in Fig. 5 and the error increases with $H_{\mathrm{ay}}$.
@@
In an equal-free-energy contour including the saddle point ($\theta_{\mathrm{oct}}=\pi/2$, $\varphi_{\mathrm{oct}}=0$),
@@
the relationship between $m_{\mathrm{oct},z}$ and $\varphi_{\mathrm{oct}}$ is given as
@@
×$\times\int_{\pi/2}^{0}\alpha(-3M_{\mathrm{s}}H_{\mathrm{K}}\sin\varphi_{\mathrm{oct}}\cos\varphi_{\mathrm{oct}}-3M_{\mathrm{s}}H_{\mathrm{x,eff}}\sin\varphi_{\mathrm{oct}})$
×$\times\sqrt{\frac{H_{\mathrm{J}}}{H_{\mathrm{K}}\sin^{2}\varphi_{\mathrm{oct}}+2H_{\mathrm{x,eff}}(1-\cos\varphi_{\mathrm{oct}})}}(2H_{\mathrm{K}}\sin\varphi_{\mathrm{oct}}\cos\varphi_{\mathrm{oct}}+2H_{\mathrm{x,eff}}\sin\varphi_{\mathrm{oct}})\mathrm{d}\varphi_{\mathrm{oct}}\Bigg{|}.$
Numerically integrating the above equation allows us to compute $\quantity[A\quantity(\frac{\delta E}{k_{\mathrm{B}}T})]^{-1}$, which can again be approximated to $1$ with an error of less than $5\%$ for $\Delta E_{0}\geq 2k_{\mathrm{B}}T$, as shown in Fig. 6(a) and the error exponentially decreases as $\Delta E_{0}$ increases.
@@
×$\times\int_{\pi/2}^{0}\alpha(-3M_{\mathrm{s}}H_{\mathrm{K}}\sin\varphi_{\mathrm{oct}}\cos\varphi_{\mathrm{oct}}-3M_{\mathrm{s}}H_{\mathrm{x,eff}}\sin\varphi_{\mathrm{oct}})$
×$\times\sqrt{\frac{H_{\mathrm{J}}}{H_{\mathrm{K}}\sin^{2}\varphi_{\mathrm{oct}}-2H_{\mathrm{x,eff}}(1+\cos\varphi_{\mathrm{oct}})}}(2H_{\mathrm{K}}\sin\varphi_{\mathrm{oct}}\cos\varphi_{\mathrm{oct}}+2H_{\mathrm{x,eff}}\sin\varphi_{\mathrm{oct}})\mathrm{d}\varphi_{\mathrm{oct}}\Bigg{|}.$
Even in this case $\quantity[A\quantity(\frac{\delta E}{k_{\mathrm{B}}T})]^{-1}\approx 1$ with an error of less than $2\%$ when $\Delta E_{0}\geq 2k_{\mathrm{B}}T$, as shown in Fig. 6(b).
.3 Numerical Simulation
.3.1 Escape time definition
@@
  • $\mathbf{m}_{\mathrm{oct}}$ switching from up state to down state is assumed to occur when $\mathbf{m}_{\mathrm{oct,y}}$ crosses $-\xi$,
@@
.3.2 Spin configuration
The unit cell of $\mathrm{Mn_{3}Sn}$ is composed of 6 $\mathrm{Mn}$ spins in two stacked kagome planes so the complete form of the Hamiltonian consisting 6 spins is Liu and Balents (2017):
@@
We further perform numerical simulations using both the 3-spin and 6-spin configurations. As shown in Fig. 8, the escape times from the two configurations overlap, indicating that the use of the 3-spin configuration in the main text is justified.
@@
.3.3 Damping coefficient dependence
Figure 9: Field-assisted escape time with $\alpha$ ranging from $10^{-4}$ to $10^{-2}$ and $\Delta E_{0}=3k_{\mathrm{B}}T$
@@
To investigate the intermediate-to-high damping regime and validate the applicable range of our theoretical model, we conducted numerical simulations over a broad range of damping coefficients $\alpha$ from $10^{-4}$ to $10^{-2}$. As shown in Fig. 9, the theoretical predictions for $H_{\mathrm{ax}}$-assisted escape remain valid across this entire range, whereas the theory fails for $H_{\mathrm{ay}}$ when $\alpha<5\times 10^{-3}$.
.3.4 Down-to-up escape time analysis
Figure 10: The schematic of escape time from down state to up state while numerical extracted data is in markers and analytical predictions are plotted with lines. $\Delta E_{0}$ ranges from $2k_{\mathrm{B}}T$ to $5k_{\mathrm{B}}T$ and field is applied along (a) $x$-axis and (b) $y$-axis.

### v2

@@ I.1.4 Analysis of depopulation factor
$\delta\mathcal{F}_{\text{oct}}=\int_{\varphi_{\text{oct}}}{-\alpha}\left(\frac{\partial\mathcal{F}_{\text{oct}}}{\partial m_{\mathrm{oct},z}}\right)\,d\varphi_{\text{oct}}+\int_{m_{\mathrm{oct},z}}{\alpha}\left(\frac{\partial\mathcal{F}_{\text{oct}}}{\partial\varphi_{\text{oct}}}\right)\,dm_{\mathrm{oct},z}.$ (33)
In order to calculate the integration numerically, we have to get the $m_{\mathrm{oct,z}}-\varphi_{\mathrm{oct}}$ relationship and make the integration $\varphi_{\mathrm{oct}}$-dependent only.
  • Depopulation factor with $H_{\mathrm{ay}}$
@@
and, therefore, the $m_{\mathrm{oct},z}-\varphi_{\mathrm{oct}}$ relationship is given as:
@@
$\times\int_{\pi/2}^{\pi}\alpha(-3M_{\mathrm{s}}H_{\mathrm{K}}\sin\varphi_{\mathrm{oct}}\cos\varphi_{\mathrm{oct}}+3M_{\mathrm{s}}H_{\mathrm{y,eff}}\cos\varphi_{\mathrm{oct}})$
$\times\sqrt{\frac{H_{\mathrm{J}}H_{\mathrm{K}}}{H_{\mathrm{K}}^{2}\sin^{2}\varphi_{\mathrm{oct}}+H_{\mathrm{y,eff}}^{2}-2H_{\mathrm{K}}H_{\mathrm{y,eff}}\sin\varphi_{\mathrm{oct}}}}\quantity(2H_{\mathrm{K}}^{2}\sin\varphi_{\mathrm{oct}}\cos\varphi_{\mathrm{oct}}-2H_{\mathrm{K}}H_{\mathrm{y,eff}}\cos\varphi_{\mathrm{oct}})d\varphi_{\mathrm{oct}}.$
Numerically integrating the above equation enables us to compute the inverse of the depopulation factor $\quantity[A\quantity(\frac{\delta E}{k_{\mathrm{B}}T})]^{-1}$. For the range of $H_{\mathrm{ay}}$ and $\mathcal{V}$ considered in this work, $\quantity[A\quantity(\frac{\delta E}{k_{\mathrm{B}}T})]^{-1}\approx 1$ with less than $1\%$ error, as presented in Fig. 5 and the error increases with $H_{\mathrm{ay}}$.
@@
Due to the asymmetry of the two saddle points induced by $H_{\mathrm{ax}}$, their free energies are no longer identical; consequently, the two equal-free-energy contours need to be treated separately. In an equal-free-energy contour including the saddle point ($\theta_{\mathrm{oct}}=\pi/2$, $\varphi_{\mathrm{oct}}=0$),
@@
the $m_{\mathrm{oct},z}-\varphi_{\mathrm{oct}}$ relationship is given as:
@@
$\times\int_{\pi/2}^{0}\alpha(-3M_{\mathrm{s}}H_{\mathrm{K}}\sin\varphi_{\mathrm{oct}}\cos\varphi_{\mathrm{oct}}-3M_{\mathrm{s}}H_{\mathrm{x,eff}}\sin\varphi_{\mathrm{oct}})$
$\times\sqrt{\frac{H_{\mathrm{J}}}{H_{\mathrm{K}}\sin^{2}\varphi_{\mathrm{oct}}+2H_{\mathrm{x,eff}}(1-\cos\varphi_{\mathrm{oct}})}}(2H_{\mathrm{K}}\sin\varphi_{\mathrm{oct}}\cos\varphi_{\mathrm{oct}}+2H_{\mathrm{x,eff}}\sin\varphi_{\mathrm{oct}})\mathrm{d}\varphi_{\mathrm{oct}}\Bigg{|}.$
Numerically integrating the above equation allows us to compute $\quantity[A\quantity(\frac{\delta E}{k_{\mathrm{B}}T})]^{-1}$, which can again be approximated to $1$ with an error of less than $1\%$ for $\Delta E_{0}\geq 3k_{\mathrm{B}}T$, as shown in Fig. 6(a) and the error exponentially decreases as $\Delta E_{0}$ increases.
@@
$\times\int_{\pi/2}^{0}\alpha(-3M_{\mathrm{s}}H_{\mathrm{K}}\sin\varphi_{\mathrm{oct}}\cos\varphi_{\mathrm{oct}}-3M_{\mathrm{s}}H_{\mathrm{x,eff}}\sin\varphi_{\mathrm{oct}})$
$\times\sqrt{\frac{H_{\mathrm{J}}}{H_{\mathrm{K}}\sin^{2}\varphi_{\mathrm{oct}}-2H_{\mathrm{x,eff}}(1+\cos\varphi_{\mathrm{oct}})}}(2H_{\mathrm{K}}\sin\varphi_{\mathrm{oct}}\cos\varphi_{\mathrm{oct}}+2H_{\mathrm{x,eff}}\sin\varphi_{\mathrm{oct}})\mathrm{d}\varphi_{\mathrm{oct}}\Bigg{|}.$
Even in the case of $H_{\mathrm{ax}}$-assisted dynamics, $\quantity[A\quantity(\frac{\delta E}{k_{\mathrm{B}}T})]^{-1}\approx 1$ with an error of less than $1\%$ when $\Delta E_{0}\geq 3k_{\mathrm{B}}T$, as shown in Fig. 6(b).
I.2 Numerical simulation of escape time
I.2.1 Details and benchmarking of LLG
The numerical simulation results presented in the main text utilize our in-house developed solver on the CUDA platform. The solver is used to run more than 1000 simulations allows us to collect statistically reliable data for analyzing the effect of thermal noise on the octupole dynamics. The LLG equation for sublattice $i=1,2,3$ is given as
$\partialderivative{\mathbf{m}_{i}}{t}=-\gamma(\mathbf{m}_{i}\crossproduct\mathbf{H}_{i}^{\mathrm{eff}})+\alpha\quantity(\mathbf{m}_{i}\times\partialderivative{\mathbf{m}_{i}}{t}),$ (43)
where $\mathbf{H}_{i}^{\mathrm{eff}}$, the effective field for sublattice $i$, is defined as
$\mathbf{H}_{i}^{\mathrm{eff}}=-\frac{1}{M_{\mathrm{s}}}\frac{\partial\mathcal{H}(\mathbf{m})}{\partial\mathbf{m}_{i}}+\mathbf{H}_{\mathrm{thermal},i}.$ (44)
The effective field includes internal fields due to exchange and DM interactions between sublattices, field due to uniaxial anisotropy along respective easy axes, externally applied magnetic field, and the thermal field $\mathbf{H}_{\mathrm{thermal},i}=\sqrt{\frac{2\alpha k_{\mathrm{B}}T}{\gamma\mu_{0}M_{\mathrm{s}}^{2}\mathcal{V}\Delta t}}\bm{\xi}^{i}_{t}$, where $\bm{\xi}^{i}_{t}\sim\mathcal{N}(0,1)$. Here, $\gamma$ and $\mu_{0}$ denote the gyromagnetic ratio and vacuum permeability, respectively, while $\Delta t=10^{-15}~\mathrm{s}$ is the time step used in the simulation. The LLG equation of each sublattice is numerically solved with finite different Heun method Ament et al. (2016).
To verify the results of our numerical implementation, firstly, we benchmarked our extracted data (in the absence of thermal noise) against previously published results Shukla, Qian, and Rakheja (2024) as depicted in Fig. 7(a), Secondly, we benchmarked our extracted numerical data (with thermal noise) against Eq. (7) of Ref. [44], as depicted in Fig. 7(b) by symbols and line, respectively. We then proceeded to investigate the field-driven dynamics in single-domain Mn3Sn in the presence of thermal noise.
Figure 7: Comparison of our in-house LLG solver against published analytical predictions. (a) Comparison against field-assisted SOT-driven dynamics without thermal noise Shukla, Qian, and Rakheja (2024). The white lines are predicted threshold current density between non-switching ($\mathrm{I}$), switching ($\mathrm{II}$) and rotation ($\mathrm{III}$). (b) Comparison against field-free thermal fluctuations Konakanchi et al. (2025). The symbols represent the numerically simulated escape time without external field and the solid line is the theoretical prediction.
I.2.2 Extraction of escape time from LLG
@@
  • $\mathbf{m}_{\mathrm{oct}}$ switching from up state to down state is assumed to occur when $\mathbf{m}_{\mathrm{oct,y}}$ crosses $\xi$,
@@
I.2.3 Two Kagome-plane versus one Kagome-plane model of spin configuration
The unit cell of $\mathrm{Mn_{3}Sn}$ is composed of 6 $\mathrm{Mn}$ spins in two stacked Kagome planes but the one-Kagome-plane is also widely adopted in theoretical and numerical analysis Shukla and Rakheja (2022); He and Liu (2024) and it has the same accuracy within the tolerance of error for the following reasons.First of all, the order parameter itself is intra-layer, which means $\mathbf{m}_{\mathrm{oct}}$ is built from one triangle of Mn spins in one Kagome plane and this already gives the correct symmetry that underlines the large Berry curvature Suzuki et al. (2017); Chen et al. (2023). Secondly, the ABAB stacked Kagome planes share the same octupolar chirality so most in-plane transport and switching results does not qualitatively change between one-plane and two-plane models. Thirdly, the inter-layer interactions is modest compared with intra-layer interactions. To validate for this, we conduct a comparison through numerical simulations. The complete form of the Hamiltonian consisting 6 spins is Liu and Balents (2017):
@@ I.2.3 Two Kagome-plane versus one Kagome-plane model of spin configuration
We perform coupled LLG equation simulations using Eq. 45. As shown in Fig. 9, the escape times from the two-Kagome-plane model overlap with those from the one-Kagome-plane model used in the main text. These numerical results confirm that the one-Kagome-plane model is sufficient for investigating the thermal dynamics of Mn3Sn.
@@
I.3 Additional analysis of escape time
I.3.1 Differences between numerical and analytical models of escape time
As depicted in Fig. (3) of the manuscript, theoretical predictions of Eq. (5)-(8) deviate from the numerical simulation results at high magnetic fields. To analyze this deviation, we rewrite Eq. (5) in the form $\tau_{\mathrm{esc}}=\tau_{0}\exp(\Delta E/k_{\mathrm{B}}T)$ and analyze $\tau_{0}$ and $\exp(\Delta E/k_{\mathrm{B}}T)$ separately. For the $H_{\mathrm{ax}}$-assisted up-to-down escape, the dominant pathway is the escape from $\varphi_{\mathrm{oct}}=\pi$, and we therefore focus on this process. As shown in Fig. 10, $\tau_{0}$ and $\exp(\Delta E_{x}/k_{\mathrm{B}}T)$ exhibit opposite dependencies on $H_{\mathrm{ax}}$. With increasing field, the reduction in the exponential factor may not compensate for the growth in $\tau_{0}$, leading to an increase in the predicted $\tau_{\mathrm{esc}}$. This trend is more pronounced in the case of $H_{\mathrm{ay}}$-assisted escape (Fig. 11), where $\tau_{0}$ grows more rapidly when the field is applied along the y-y axis. This explains the abnormal increase in escape time with increasing field strength, which underlies the deviation of the theoretical predictions from numerical simulations. This deviation warrants further investigation, as the applicability of the perturbation theory under increasing external field strength remains uncertain. Both the linearization of the LLG equation and the eigenvalue analysis of the Hessian matrix rely on a Hamiltonian that depends on $\mathbf{m}_{\mathrm{oct}}$, which is itself derived from perturbative assumptions. If a strong external field disrupts the rigid 120 spin configuration characteristic of Mn3Sn, the perturbative expression in Eq. (2) may significantly deviate from the full Hamiltonian in Eq. (1). Consequently, this deviation can lead to larger errors in the estimation of both the attempt time $\tau_{0}$ and the energy barrier $\Delta E$.
[…]
I.3.2 Sensitivity of escape time to variations in material parameters
The up-to-down escape time is analytically derived from Eqs. (3)–(8) in the main text, making it difficult to directly assess its dependence on the magnetic parameters listed in Table I of the main text. To address this, we plot in Fig. 12 a heat map of the analytically obtained up-to-down escape time by sweeping the external field and other magnetic parameters, while fixing $\Delta E_{0}=4~k_{\mathrm{B}}T$ in Fig. 12 and fixing $\mathcal{V}=80\times 80\times 8~\mathrm{nm^{3}}$ in Fig. 13.
Figure 12: The up-to-down escape time with $\Delta E_{0}=4k_{\mathrm{B}}T$ by sweeping external field and magnetic paramters (a)&(b) $J_{\mathrm{E}}$, (c)&(d) $D_{\mathrm{M}}$, (e)&(f) $K_{\mathrm{u}}$, (g)&(h) $\delta_{\mathrm{E}}$, (i)&(j) $M_{\mathrm{s}}$, in Table $\mathrm{I}$ in the main text. The left column corresponds to $H_{\mathrm{ax}}$, while the right column corresponds to $H_{\mathrm{ay}}$.
Figure 13: The up-to-down escape time with $\mathcal{V}=80\times 80\times 8~\mathrm{nm^{3}}$ by sweeping external field and magnetic paramters (a)&(b) $J_{\mathrm{E}}$, (c)&(d) $D_{\mathrm{M}}$, (e)&(f) $K_{\mathrm{u}}$, (g)&(h) $\delta_{\mathrm{E}}$, (i)&(j) $M_{\mathrm{s}}$, in Table $\mathrm{I}$ in the main text. The left column corresponds to $H_{\mathrm{ax}}$, while the right column corresponds to $H_{\mathrm{ay}}$.
$H_{\mathrm{ax}}$-assisted and $H_{\mathrm{ay}}$-assisted thermal escape time exhibit qualitatively similar dependence trends under identical parameter variations, while $\tau_{\mathrm{esc}}$ is quantitatively more sensitive in the $H_{\mathrm{ay}}$-assisted case.
As shown in Figs. 12(a)&(b), increasing $J_{\mathrm{E}}$ reduces $\tau_{\mathrm{esc}}$ in the low-field region, indicating that the strong exchange interaction governs the ultrafast thermal fluctuation frequency of Mn3Sn. At higher fields, however, the dependence of $\tau_{\mathrm{esc}}$ on $J_{\mathrm{E}}$ reverses, since the attempt frequency in Eq. (7) is not monotonically dependent on $J_{\mathrm{E}}$. In contrast, Figs. 13(a)&(b) show that $\tau_{\mathrm{esc}}$ increases monotonically with $J_{\mathrm{E}}$, which can be attributed to its contribution to the energy barrier through $H_{\mathrm{K}}$.In contrast, $D_{\mathrm{M}}$ exerts only a limited influence on $\tau_{\mathrm{esc}}$, since it is about an order of magnitude smaller than $J_{\mathrm{E}}$; consequently, the effect of $D_{\mathrm{M}}$ is suppressed by $J_{\mathrm{E}}$.
According to Eq. (7), the attempt frequency is positively related to the magnitude of the uniaxial anisotropy $H_{\mathrm{K}}$, such that, when $\Delta E_{0}$ is fixed, the escape time decreases with increasing $H_{\mathrm{K}}$. Since $H_{\mathrm{K}}$ is proportional to both $K_{\mathrm{u}}$ and $\delta_{\mathrm{E}}$, larger values of these parameters accelerate the dynamics, as shown in Fig. 12(e)–(h). Although a larger $K_{\mathrm{u}}$ corresponds to a stronger uniaxial field that favors alignment of the sublattices along their easy axes, maintaining a constant $\Delta E_{0}$ requires a reduction in size. This reduction enhances the thermal field, thereby further accelerating the dynamics. In contrast, $M_{\mathrm{s}}$ exerts the opposite influence: both $H_{\mathrm{K}}$ and the thermal field decrease with increasing $M_{\mathrm{s}}$, leading to longer escape times for larger $M_{\mathrm{s}}$.
Despite the positive dependence of $K_{\mathrm{u}}$ and $|\delta_{\mathrm{E}}|$ on the attempt frequency, Figs. 13(e)&(h) show that $\tau_{\mathrm{esc}}$ increases with these parameters. This is because, under the fixed-volume condition, $\tau_{\mathrm{esc}}$ is predominantly governed by $\Delta E$. Eqs. (3)&(4) clearly indicate that $\Delta E$ is positively dependent on $H_{\mathrm{K}}$, and therefore also on $K_{\mathrm{u}}$ and $|\delta_{\mathrm{E}}|$. In contrast, $M_{\mathrm{s}}$ is positively dependent on $\Delta E$ despite its negative dependence on $H_{\mathrm{K}}$, which explains why $\tau_{\mathrm{esc}}$ in Figs. 13(i)&(j) increases with $M_{\mathrm{s}}$.
I.3.3 Impact of damping coefficient
@@
To investigate the intermediate-to-high damping regime and validate the applicable range of our theoretical model, we conducted numerical simulations over a broad range of damping coefficients $\alpha$ from $10^{-4}$ to $10^{-2}$. As shown in Fig. 14, the theoretical predictions for $H_{\mathrm{ax}}$-assisted escape remain valid across this entire range, whereas the theory fails for $H_{\mathrm{ay}}$ when $\alpha<5\times 10^{-3}$. As $\alpha$ decreases below $5\times 10^{-3}$, $\tau_{\mathrm{escy,\uparrow}}$ for $\alpha=10^{-4}$ becomes even larger than that for $\alpha=5\times 10^{-4}$. The mechanism behind this behavior remains unclear, but one possible explanation is that, similar to the case of superparamagnets, energy exchange with the thermal bath becomes slow Kramers (1940), while precessions along the effective field increase Brown Jr (1963); both effects act to delay the escape process.
Figure 14: Field-assisted escape time with $\alpha$ ranging from $10^{-4}$ to $10^{-2}$ and $\Delta E_{0}=4k_{\mathrm{B}}T.$
I.3.4 Escape time for down-to-up dynamics
@@
Figure 15: The schematic of escape time from down state to up state while numerical extracted data is in markers and analytical predictions are plotted with lines. $\Delta E_{0}$ ranges from $3k_{\mathrm{B}}T$ to $5k_{\mathrm{B}}T$ and field is applied along (a) $x$-axis and (b) $y$-axis.